OpenStudy (anonymous):

r = 6/(sin(theta) + cos(theta))

OpenStudy (anonymous):

\[r=\frac{6}{\sin\theta + \cos\theta}\] What exactly do you need help with here?

OpenStudy (anonymous):

convert from polar form to regular form

OpenStudy (anonymous):

Okay, so you know that \[x=r\cos\theta\\ y=r\sin\theta\\ \tan\theta=\frac{y}{x}, \text{ right?}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

And one more conversion: \[r=\sqrt{x^2+y^2}.\] Okay, so in the denominator, you can rewrite the original equation as \[\sqrt{x^2+y^2}=\frac{6}{\sin{(\tan^{-1}\frac{y}{x}})+\cos{(\tan^{-1}\frac{y}{x}})}\] Do you think you can solve that?

OpenStudy (anonymous):

No, I don't think I can because I have tried several times and haven't been successful

OpenStudy (anonymous):

Okay, do you know how to rewrite the terms in the denominator?

OpenStudy (anonymous):

nope, I haven't seen anything like that before

OpenStudy (anonymous):

Alright then. Take a look at this picture: If you need any help understanding it, let me know.